In a certain population an average of 11 new cases of esophageal cancer are diagnosed each year. If the annual incidence of esophageal cancer follows a Poisson distribution, find the probability that in a given year the number of newly diagnosed cases of esophageal cancer will be at least 2.
Answer #1

Answer:

Given,

To determine the probability that in a given year the number of newly diagnosed cases of esophageal cancer will be at least 2

\lambda = 11

P(X >= 2) = 1 - P(X <= 1)

= 1 - [P(X = 0) + P(X = 1)]

= 1 - [e^-11*11^0 / 0! + e^-11*11^1 / 1!]

= 1 - [0.0000167 + 0.000184]

= 1 - 0.0002007

= 0.9998

P(X >= 2) = 0.9998